# Ab uno disce omnes: Single-harmonic search for extreme mass-ratio inspirals

### v1

@@ Abstract
Extreme mass-ratio inspirals (EMRIs) are one of the key sources of gravitational waves for space-based detectors such as LISA. However, their detection remains a major data analysis challenge due to the signals’ complexity and length. We present a semi-coherent, time-frequency search strategy for detecting EMRI harmonics without relying on full waveform templates. We perform an injection and search campaign of single mildly-eccentric equatorial EMRIs in stationary Gaussian noise. The detection statistic is constructed solely from the EMRI frequency evolution, which is modeled phenomenologically using a Singular Value Decomposition basis. The pipeline and the detection statistic are implemented in time-frequency, enabling efficient searches over one year of data in approximately one hour on a single GPU. The search pipeline achieves 94% detection probability at $\mathrm{SNR}=30$ for a false-alarm probability of $10^{-2}$, recovering the frequency evolution of the dominant harmonic to 1% relative error. By mapping the EMRI parameters consistent with the recovered frequency evolution, we show that the semi-coherent detection statistic enables a sub-percent precision estimation of the EMRI intrinsic parameters. These results establish a computationally efficient framework for constructing EMRI proposals for the LISA global fit.
@@ I Introduction
Extreme mass-ratio inspirals (EMRIs) involve the capture and inspiral of a compact stellar remnant into a massive black hole (MBH). These systems are one of the key gravitational-wave (GW) sources of future space-based detectors like the Laser Interferometer Space Antenna (LISA) [1], TianQin [2], and Taiji [3]. The scientific potential of EMRI observations is vast. They offer a unique perspective of the properties and environments of MBHs in galactic cores [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. They also provide an effective means of constraining fundamental physics, being both highly sensitive to effects predicted by modified gravity theories [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] and effective probes of cosmological expansion  [31, 32, 33]. Despite this strong case for EMRI science, identifying their signals in noisy data is still a largely unsolved data analysis problem due to several factors:
@@ II.3 Singular Value Decomposition basis for EMRI frequency evolution
We checked that such a prior covers most of the frequency tracks in the EMRI parameter space of Table 1.
@@ II.5 Simulation based inference for the EMRI inverse problem
However, since the recovered frequency evolution is subject to uncertainties and multiple EMRI parameters can give rise to very similar frequency tracks [45]. We must account for these factors in our estimation process. This can be achieved by incorporating a Bayesian framework, where we treat the EMRI parameters as random variables and update our beliefs about their values based on the observed data. To do so we need to estimate the uncertainty on the recovered frequency evolution. This can be done by analyzing the distribution of recovered frequency evolutions from multiple searches. If we have an estimate of the uncertainty $\sigma$ with which our pipeline recovers on average the frequency evolution, we can define a data generation process $f_{\alpha}^{\rm rec}=f_{\alpha}(\theta)+\epsilon_{\alpha}$, where $\epsilon_{\alpha}$ is a Gaussian noise term with zero mean and standard deviation $\sigma$. The noise characteristics in the recovered frequency evolution are generally unknown and may deviate from Gaussianity. Moreover, the recovered signal may not align perfectly with the EMRI model, as it is obtained via a phenomenological SVD basis fit. Nonetheless, assuming Gaussian noise provides a practical starting point for the inverse problem and remains valid as long as the errors are sufficiently large to avoid biasing the mapping.
@@ III.4 EMRI parameter identification
We found that the highest detection-statistic values were concentrated in a narrow slice of the parameter space (Fig. 7), with a gentle slope toward lower primary masses and a steeper decline toward higher primary masses. The maximum sample has a relative differences of 1.5% for $\log_{10}m_{1}$ and of order 0.1-0.03% for $m_{2},a,e_{f},T_{\rm pl}$. The SBI proposal seems to have guided us along the right correlations of EMRI parameters, but it remains to be seen whether the number of samples used is sufficient to establish a robust global maximum of the detection statistic across different injections.
@@ IV.2 Future improvements
  • Multi-harmonic extension. Including multiple harmonics would improve sensitivity for eccentric systems and provide additional parameter constraints. The time-frequency framework naturally accommodates multi-mode searches with proportional computational-cost scaling.
@@ IV.3 Astrophysical implications
For reference, we show the primary and secondary masses (median) and redshifts of current LIGO/Virgo/KAGRA (LVK) detections [102, 103, 104, 105, 106, 107, 108, 109] as red dots in the left panel. Our systems extend to higher redshifts and lower secondary masses. However, future analyses will need to expand the prior range to include larger masses. Even though EMRIs and LVK sources originate from distinct formation channels, the secondary object in an EMRI may still form through stellar evolution before being captured. This offers a complementary perspective on the population of stellar-mass compact objects in galactic centers. The secondary-mass distribution highlights the detectability of lighter stellar remnants ($m_{2}\lesssim 30\,M_{\odot}$) out to $z\sim 1$, offering a direct probe of the demographics of stellar-mass black holes and neutron stars in galactic nuclei close to cosmic noon—the epoch of peak star formation and galaxy merger activity around $z\sim 1$–2 [111].
EMRIs formed through disk-assisted channels [112], sometimes called “wet EMRIs,” are expected to occur in a fraction $\sim 10^{-2}$$10^{-1}$ of galactic nuclei. In these systems, interactions between the compact object and the surrounding accretion disk can lead to high-energy electromagnetic emission, making them promising candidates for multi-messenger observations. Two particularly interesting class of such events are the so-called quasi-periodic eruptions (QPEs) and quasi-periodic oscillations (QPOs) The former are semi-regular bursts observed in soft X-ray bands and the latter are less abrupt events due to oscillations in the X-ray flux of active galactic nuclei (AGNs). Under the assumption that these phenomena are powered by EMRIs, we show in the right panel of Fig. 8 the massive black hole masses and redshifts of QPEs and QPOs [113, 114, 115, 116, 117, 118, 119, 120, 121] from Table I of [110]. The EMRI sources considered in this work are detected at higher redshifts than these events, showing the potential of EMRIs to explore the multi-messenger opportunities of QPEs and QPOs. However, our prior does not include higher primary masses so future works will need to assess the detectability of EMRIs with large primary masses $\sim 10^{7}M_{\odot}$, where the impact of overlapping galactic binaries might play a crucial role.
@@
In contrast, EMRIs formed via the two-body relaxation channel in nuclear star clusters are expected to be highly eccentric and will likely require a dedicated search strategy [122]. Assessing the minimum SNR required for a confident detection based on a single harmonic would be an important next step toward characterizing this pipeline’s performance in the high eccentricity regime.
@@ IV.4 Implications for the LISA global fit
In the context of the LISA global fit, the harmonic finder introduced in this work plays a key role in identifying EMRI candidates. Combined with the SBI framework, the search strategy provides promising initial conditions for full-waveform template placement. However, our analysis indicated that additional refinements are necessary to locate the true maximum of the detection statistic across the parameter space. The search and identification pipelines developed here do not rely on waveform templates; instead, they exploit only the relationship between EMRI parameters and their frequency evolution. This approach significantly reduces computational cost and enables scalable candidate identification across a broad parameter space. However, further investigation is required to assess whether the statistic defined in Eq. (6) is sufficient for robust EMRI identification throughout the full parameter space.
Moreover, the long-duration nature of EMRI signals makes them particularly susceptible to non-stationary noise and instrumental artifacts [124, 125]. Recently, Ref. [126] showed that for stellar origin black hole binaries the framework presented here is robust against non-stationary noise. Understanding how these factors impact EMRI detection and parameter estimation will be crucial. Our time-frequency framework lays the groundwork for addressing these challenges.
Galactic binaries and massive black hole binaries will influence EMRI detection in distinct ways due to their differing characteristics [84]. The structure of our search pipeline enables the use of detected frequency evolution to mask the data, potentially isolating EMRIs for targeted follow-up. Exploring these interactions will be essential for developing a dedicated EMRI module within the LISA global fit framework.
@@ Acknowledgements.
We thank Jonathan Gair and Philippa Cole for discussions. L. S thanks Gijs Nelemans for providing the computational resources that enabled the large-scale simulations and data analysis presented in this work. L. S. is supported by the European Space Agency Research Fellowship programme. R.T. and D.G. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. C.E.A.C-B. is supported by UKSA Space Agency grant UKRI971. D.G. is supported by MUR Young Researchers Grant No. SOE2024-0000125, and MSCA Fellowship No. 101149270–ProtoBH. This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-10027. Computational work was performed at CINECA with allocations through INFN and the University of Milano-Bicocca, and at NVIDIA with allocations through the Academic Grant program. The authors thank participants and organizers of the Scientific Machine Learning for Gravitational Wave Astronomy workshop held at ICERM (Providence, RI), which is supported by NSF Grant No. DMS-1929284. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This work makes use of the Black Hole Perturbation Toolkit, few [42, 43, 44, 35], numpy [127], matplotlib [128], scipy [94], jax [66], optax [73], interpax [129], and pastamarkers [130, 131].

### v2

@@ Abstract
Extreme mass-ratio inspirals (EMRIs) are one of the key sources of gravitational waves for space-based detectors such as LISA. However, their detection remains a major data analysis challenge due to the signals’ complexity and length. We present a semi-coherent, time-frequency search strategy for detecting EMRI harmonics without relying on full waveform templates. We perform an injection and search campaign of single mildly-eccentric equatorial EMRIs in stationary Gaussian noise. The detection statistic is constructed solely from the EMRI frequency evolution, which is modeled phenomenologically using a Singular Value Decomposition basis. The pipeline and the detection statistic are implemented in time-frequency, enabling efficient searches over one year of data in approximately one hour on a single Graphic Processing Unit (GPU). The search pipeline achieves 94% detection probability at $\mathrm{SNR}=30$ for a false-alarm probability of $10^{-2}$, recovering the frequency evolution of the dominant harmonic within 1% relative error. By mapping the EMRI parameters consistent with the recovered frequency evolution, we show that the semi-coherent detection statistic enables a sub-percent precision estimation of the EMRI intrinsic parameters. These results establish a computationally efficient framework for constructing EMRI proposals for the LISA global fit.
@@ I Introduction
Extreme mass-ratio inspirals (EMRIs) involve the capture and inspiral of a compact stellar remnant into a massive black hole (MBH). These systems are one of the key gravitational-wave (GW) sources of future space-based detectors like the Laser Interferometer Space Antenna (LISA) [44], TianQin [85], and Taiji [62]. The scientific potential of EMRI observations is vast. They offer a unique perspective of the properties and environments of MBHs in galactic cores [106, 130, 76, 50, 25, 72, 74, 104, 56, 53]. They also provide an effective means of constraining fundamental physics, being both highly sensitive to effects predicted by modified gravity theories [133, 134, 107, 14, 96, 26, 28, 24, 39, 17, 36, 29, 55, 120, 105, 23, 102] and effective probes of cosmological expansion  [84, 118, 79]. Despite this strong case for EMRI science, identifying their signals in noisy data remains a largely unsolved data analysis problem due to several factors:
@@ II.3 Singular Value Decomposition basis for EMRI frequency evolution
We verified that such a prior covers most of the frequency tracks in the EMRI parameter space of Table 1.
@@ II.5 Simulation based inference for the EMRI inverse problem
However, since the recovered frequency evolution is subject to uncertainties, multiple EMRI parameters can give rise to very similar frequency tracks [38]. We must account for these factors in our estimation process. This can be achieved by incorporating a Bayesian framework, where we treat the EMRI parameters as random variables and update our beliefs about their values based on the observed data. To do so we need to estimate the uncertainty on the recovered frequency evolution. This can be done by analyzing the distribution of recovered frequency evolutions from multiple searches. If we have an estimate of the uncertainty $\sigma$ with which our pipeline recovers on average the frequency evolution, we can define a data generation process $f_{\alpha}^{\rm rec}=f_{\alpha}(\theta)(1+\epsilon_{\alpha})$, where $\epsilon_{\alpha}$ is a Gaussian noise term with zero mean and standard deviation $\sigma$. The noise characteristics in the recovered frequency evolution are generally unknown and may deviate from Gaussianity. Moreover, the recovered signal may not align perfectly with the EMRI model, as it is obtained via a phenomenological SVD basis fit. Nonetheless, assuming Gaussian noise provides a practical starting point for the inverse problem and remains valid as long as the errors are sufficiently large to avoid biasing the mapping.
@@ III.4 EMRI parameter identification
We found that the highest detection-statistic values were concentrated in a narrow slice of the parameter space (Fig. 7), with a gentle slope toward lower primary masses and a steeper decline toward higher primary masses. The maximum sample has a relative difference of 1.5% for $\log_{10}m_{1}$ and of order 0.1-0.03% for $m_{2},a,e_{f},T_{\rm pl}$. The SBI proposal seems to have guided us along the right correlations of EMRI parameters, but it remains to be seen whether the number of samples used is sufficient to establish a robust global maximum of the detection statistic across different injections.
@@ IV.2 Future improvements
  • Multi-harmonic extension. Including multiple harmonics would improve sensitivity for eccentric systems and provide additional parameter constraints. The time-frequency framework naturally accommodates multi-mode searches with proportional computational-cost scaling. Possible approaches to include higher harmonics could follow Refs. [122, 101].
@@ IV.3 Astrophysical implications
For reference, we show the primary and secondary masses (median) and redshifts of current LIGO/Virgo/KAGRA (LVK) detections [3, 4, 5, 7, 1, 8, 6, 2] as red dots in the left panel. EMRI detectable systems extend to higher redshifts than LVK sources for secondary masses smaller than $20\,M_{\odot}$. Although EMRIs and LVK sources originate from distinct formation channels, the secondary object in an EMRI may still form through stellar evolution before capture, providing a complementary perspective on stellar-mass compact object populations in galactic centers. In particular, the secondary-mass distribution reveals detectability of lighter remnants ($m_{2}\lesssim 30\,M_{\odot}$) out to $z\sim 1.5$, probing compact object demographics in galactic nuclei near cosmic noon—the epoch of peak star formation and galaxy merger activity around $z\sim 1$–2 [51]. We also notice that our prior range is limited to source-frame secondary masses of $80\,M_{\odot}$, and future analyses should be extending this range.
EMRIs formed through disk-assisted channels [93], sometimes called “wet EMRIs,” are expected to occur in a fraction $\sim 10^{-2}$$10^{-1}$ of galactic nuclei. In these systems, interactions between the compact object and the surrounding accretion disk can lead to high-energy electromagnetic emission, making them promising candidates for multi-messenger observations. Two interesting classes of such events are the so-called quasi-periodic eruptions (QPEs) and quasi-periodic oscillations (QPOs). The former are semi-regular bursts observed in soft X-ray bands and the latter are less abrupt events due to oscillations in the X-ray flux of active galactic nuclei (AGNs). Under the assumption that these phenomena are powered by EMRIs, we show in the right panel of Fig. 8 the massive black hole masses and redshifts of QPEs and QPOs [90, 11, 58, 57, 10, 81, 94, 95, 60] from Table I of [73]. The EMRI sources considered in this work are detected at higher redshifts than these QPE and QPO events [73]. Future work will need to investigate the sky localization precision of follow-up EMRI gravitational-wave analyses required to enable electromagnetic observations for multi-messenger studies of these systems. We notice that our prior does not include source-frame primary masses larger than $5\times 10^{6}\,M_{\odot}$ so future works will need to assess the detectability of EMRIs with large primary masses $\sim 10^{7}M_{\odot}$, where the impact of overlapping galactic binaries might play a crucial role.
@@
In contrast, EMRIs formed via the two-body relaxation channel in nuclear star clusters are expected to be highly eccentric and will likely require a dedicated search strategy [87, 9]. Assessing the minimum SNR required for a confident detection based on a single harmonic would be an important next step toward characterizing this pipeline’s performance in the high eccentricity regime.
@@ IV.4 Implications for the LISA global fit
In the context of the LISA global fit, the pipeline introduced in this work can identify EMRI candidates and, combined with the SBI framework, provide constrained parameter space regions for follow-up analyses using full-waveform templates. Our analysis showed an example of such a follow-up using the statistic defined in Eq. (6). The search and identification pipelines developed here do not rely on waveform templates; instead, they exploit only the relationship between EMRI parameters and their frequency evolution. This approach significantly reduces computational cost and enables scalable candidate identification across a broad parameter space. Further investigation is required to assess whether the statistic defined in Eq. (6) is sufficient for robust EMRI identification throughout the full parameter space.
Moreover, the long-duration nature of EMRI signals makes them more likely to be affected by non-stationary noise and instrumental artifacts [12, 35]. Recently, Ref. [22] showed that for stellar origin black hole binaries the detection statistic of Eq. (6) is robust against non-stationary noise. Understanding how these effects impact EMRI detection and parameter estimation will be crucial.
Galactic binaries and massive black hole binaries will influence EMRI detection differently based on their distinct characteristics in the time-frequency plane [75]. Future work should investigate whether the detected frequency tracks could be used to mask signals and enable targeted EMRI follow-up. Characterizing and leveraging these multi-source interactions will be crucial for integrating a robust EMRI search capability into the LISA global fit.
@@ Acknowledgements.
We thank Jonathan Gair, Elena Maria Rossi and Philippa Cole for discussions. L. S. thanks Gijs Nelemans for providing the computational resources that enabled the large-scale simulations and data analysis presented in this work. L. S. is supported by the European Space Agency Research Fellowship programme. R.T. and D.G. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. C.E.A.C-B. is supported by UKSA Space Agency grant UKRI971. D.G. is supported by MUR Young Researchers Grant No. SOE2024-0000125, and MSCA Fellowship No. 101149270–ProtoBH. This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-10027. Computational work was performed at CINECA with allocations through INFN and the University of Milano-Bicocca, and at NVIDIA with allocations through the Academic Grant program. The authors thank participants and organizers of the Scientific Machine Learning for Gravitational Wave Astronomy workshop held at ICERM (Providence, RI), which is supported by NSF Grant No. DMS-1929284. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This work makes use of the Black Hole Perturbation Toolkit, few [40, 70, 110, 37], numpy [61], matplotlib [64], scipy [121], jax [32], optax [19], interpax [45], and pastamarkers [99, 31].